MikesArcade.com Coin-op Video Arcade Technical Information Archive
  Home  |  Repair Logs  |  Game Manuals  |  Monitor Manuals  |  Online Store  |  Links  |  Contact Us
Technical Info
  Repair Logs
  Pacman fixes
  Manuals & Schematics
  Monitor Manuals
  Nintendo Game List
  Nintendo Joysticks
  Nintendo Buttons
  Data I/O Promlink 6.10
Spies Wiretap Archives
  Game Conversions
  The Mixed Bag
  PCB Pinouts and DIPs
PDF Game Artwork
  Coin door inserts, etc.
  Instructions cards
  Operation Sheets
  Control Panel & Bezels
Jukebox
  Title Strip Creator
Tutorials
  Component markings...
  Crating a game
  Build your own cabinet
Board Upgrades
  Double Donkey Kong
Wanted
  Items I'm looking for
Miscellaneous
  Links
  Contact Us
MAMETM & MAMEUI
  Visual C Project files
Certifications
  Valid HTML 4.01!
 

Repair Logs


Mario Bros (Nintendo) PCB repairs

10/10/2003

Tools: Schematics, Logic Probe

Problem: Walk sound distorted

Both Mario and Luigi walk sounds were messed up. Sounded more like a gush of air than the normal walking sound. Locating the Walk 1 and Walk 2 circuits on the schematics, it shows both LS629s at 1J and 2J do most of the work. Replacing both of these chips fixed the walk sound for both players.

Solution: Replaced bad 74LS629 chips at 1J and 2J on the CPU board


10/10/2003

Problem: All background graphics shifted up with the score barely showing on the bottom of the screen.

My first though was this was a Video board problem. I probed around the LS86 and LS377 and noticed that if I shorted pins 5 and 6 of the LS86 at 1D, that the picture would shift down to close to where it should be. Pin 6 is an input from the FLIP circuit and was stuck low, if I pulled this pin high with +5, the background would shift into place. Tracing back from pin 6 on 1D, All the chips were acting as they should. I moved then to the CPU board side of the FLIP circuit. On the CPU board, FLIP is connected first to pin 12 of the LS04 at 1H. Probing that pin, I noticed that it was LOW, I followed this trace back to the Video board where it connects to pin 9 and 10 of the LS00 at 5H. I turned off the board and check for continuity and sure enough it was open. Moving back to the CPU board, the trace leading to pin 12 of the LS04 at 1H was damaged. Repairing that trace fixed the problem.

Solutions: Pin 12 of LS04 at 1H on the CPU was out of circuit, repairing the trace shifted the screen back to normal.


< Prev Page

 

Home  |  Online Store  |  Repair Logs  |  Manuals and Schematics  |  Monitor Manuals  |  Component Markings  |  Links

All Content Copyright © 2000-2017 by MikesArcade.com